Pt 1. If a stock has a market beta greater than 1, the expected return will be less than the expected return of market portfolio.

Pt. 2 The stock of Target Corporation has a return of 36.43. The market risk premium is 16.94 percent and the risk-free rate is 8.04 percent. What is the beta on this stock? Use the CAPM Equation
Answer #1

CAPM Equation:

Rs=Rf+Beta*(Rm-Rf)

Rs=Expected Stock Return

Rf=Risk Free Rate

Rm=Return of market portfolio

Rs-Rf=Beta*(Rm-Rf)

If Beta is =1

Rs=Rm

If Beta >1 , Rs >Rm

If Beta<1, Rs<Rm

If a stock has a market beta greater than 1, the expected return will be GREATER  than the expected return of market portfolio.

Pt2

Expected Return of Stock=Rs=36.43%

Risk Free Rate =Rf=8.04%

Market Risk Premium =Rm-Rf=16.94%

36.43=8.04+Beta*16.94

(36.43-8.04)=Beta*16.94

28.39=Beta*16.94

Beta=28.39/16.94=1.6759

Beta of the stock=1.6759